40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17090     Entry time: Thu Aug 18 16:35:29 2022
Author: Cici 
Type: Update 
Category: General 
Subject: UGF linked to optical gain! 

TL;DR: When the laser has good lock, the OLTF moves up and the UGF moves over!

-----------------------------------------------------------

Figured out with Paco yesterday that when the laser is locked but kind of weakly (mirrors on the optical table sliiightly out of alignment, for example), we would get a UGF around 5 kHz, but when we had a very strong lock (adjusting the mirrors until the spot was brightest) we would get a UGF around 13-17 kHz. Attached are some plots of us going back and forth (you can kind of tell from the coherence/error that the one with the lower UGF is more weakly locked, too). Error on the plots is propagated using the coherence data (see Bendat and Piersol, Random Data, Table 9.6 for the formula). 

-------------------------------------------------------------

Want to take data next week to quantitatively compare optical gain to UGF!

Attachment 1: rpi_OLG_2022_08_17_18_03_52.pdf  36 kB  Uploaded Thu Aug 18 17:39:55 2022  | Hide | Hide all
rpi_OLG_2022_08_17_18_03_52.pdf
Attachment 2: rpi_OLG_2022_08_17_18_00_50.pdf  36 kB  Uploaded Thu Aug 18 17:40:13 2022  | Hide | Hide all
rpi_OLG_2022_08_17_18_00_50.pdf
ELOG V3.1.3-